### v1

# Exploring Time Delay Interferometry Ranging as a Practical Ranging Technique in the Bayesian Framework

Abstract
Knowledge on the inter-spacecraft optical paths (i.e. delays) is one of the key elements of time delay interferometry (TDI). Conventional method for inter-spacecraft ranging mainly relies on the pseudo-random noise (PRN) code signal modulated onto the lasers. To ensure the reliability and robustness of this ranging information, it would be highly beneficial to develop other methods which could serve as cross-validations or backups. This paper explores the practical implementation of an alternative data-driven approach - time delay interferometry ranging (TDIR) - as a ranging technique independent of the PRN signal. Distinguished from previous research, the inputs of our TDIR algorithm are interferometric measurements that have only undergone a 0.1 ms accuracy preliminary clock synchronization using the ground tracking data. This significantly relaxes the stringent requirement for clock synchronization imposed by traditional TDI procedure. In order to encode the information of time-varying light travel times and clock desynchronizations, we adopt a general polynomial parametrization to represent the delays. By framing TDIR as a Bayesian parameter estimation problem, we demonstrate our algorithm with simulated data based on the numerical orbit of Taiji. In the presence of laser frequency noise and secondary noises, the estimated median values of parameters exhibit a bias of only 5.28 ns from the “true” delays, capable of suppressing laser frequency noise to the desired level. Furthermore, we have also analysed the requirements of mitigating optical bench noise and clock noise on TDIR, and presented an illustrative example for the impact of laser locking.
preprint: APS/123-QED
I Introduction
The first detection of gravitational wave (GW) [1] marked the beginning of the gravitational-wave astronomy era. The LIGO-VIRGO-KAGRA ground-based GW detector network, whose target frequencies ranges from 1 Hz to 1 kHz, has detected over a hundred GW events originating from compact binaries. In contrast, in the upcoming decade, space-based detectors such as the Laser Interferometer Space Antenna (LISA) [2], Taiji [3, 4, 5], and Tianqin [6], aim to detect GWs in the 0.1 mHz - 1 Hz frequency band associated with enormous astrophysical and cosmological procedures [7], such as the mergers of massive black hole binaries [8], the orbiting of compact binaries within the Milky Way [9], the extreme mass ratio inspirals, and the first-order phase transition in the early universe [10], etc. As a representative example, Taiji consists of three spacecrafts (S/Cs) connected by bidirectional laser links, with a baseline armlength of 3 million kilometers. The whole constellation orbits around the sun and leads the Earth by about 20 degrees.
Albeit overcoming the limitations of the Earth’s scale and seismic noise, space-based GW detectors subject to laser frequency noise (LFN) which is coupled with the inequality of armlengths. LFN in the raw interferometric measurements can be 6 - 8 orders stronger than the targeted GW signals. To suppress this overwhelming noise source, the TDI technique was proposed to synthesise a virtual equal-arm interferometer in the post-processing stage [11, 12, 13]. Generally, the construction of TDI channels involves appropriately delaying and combining the readouts of interferometors, including the science (i.e. inter-spacecraft), reference and test-mass interferometors. According to the capability of LFN suppression, TDI combinations can be categorized into the first and second generations. For an ideal stationary unequal-arm constellation, perfect cancellation of LFN can be achieved by the first-generation combinations. While for LISA and Taiji whose armlengths vary at rates up to several m/s, it is necessary to employ the second-generation TDI combinations to account for both the inequality and time dependence of the armlengths [14].
In addition to the interferometric measurements (also known as the beatnote signals), the inter-spacecraft optical paths (i.e. delays) constitute another crucial input for TDI. Technically, the inter-spacecraft ranging can be done by modulating the PRN code onto the lasers using approximately $1\%$ of the power [15, 16, 17, 18, 19]. The resultant PRN ranging (PRNR) may suffer from several imperfections, such as ranging noise, ambiguity, modulation noise, and on-board delays, etc. TDI places stringent requirement on the precision of delays ($\mathcal{O}(10)\sim\mathcal{O}(100)\ {\rm ns}$, depending on the TDI channels [20, 21]), so the practical application of PRNR necessitates careful processing through the combination with other ranging observables [22], based on a thorough investigation and comprehensive understanding of the technical details. Due to these technical imperfections and complexities, as well as the fundamental importance of TDI for space-based GW detection, the development of other ranging methods that are independent of PRN signal and its processing step is also desirable. Such method could serve as a risk mitigation or cross verification, and hence improve the reliability and robustness of ranging.
TDIR is a data-driven approach of inter-spacecraft ranging, which does not directly measure the delays, but instead searches for the delays that minimize the residual LFN in the TDI combinations [23]. Ref. [23] proposed the basic concept of TDIR, construing it as an optimization problem aimed at minimizing the integrated power of a single TDI channel. When the delays used in TDI are close to their true values, LFN can be suppressed well below the secondary noises (i.e. the optical measurement system (OMS) noises and residual acceleration noises of the test masses). According to the principles of Bayesian statistical inference, an optimal estimator can be formulated by whitening the power of TDI channels using the noise power spectral density (PSD) of secondary noises in the Fourier domain. Researches based on this idea view TDIR as a Bayesian parameter estimation problem, which not only yields the best estimate of delays, but also outputs the confidence interval for the estimate, facilitating the assessment of ranging precision. Ref. [24] formulated the likelihood of TDIR for he first-generation Michelson channels $\{X,Y,Z\}$, ran a Markov Chain Monte Carlo (MCMC) search on the simulated data, and constrained the delays to a $\mathcal{O}(10)\ {\rm ns}$ precision. In a following up research [25], the same algorithm was extended to the second-generation TDI and time-varying delays, presuming that the delays could be parameterized by Keplerian orbital elements. No-bias estimate of the parameters was obtained on LFN-only simulations. Complementarily, a theoretical analysis on the precision of TDIR was made by Ref. [26] using the Fisher matrix formalism. In addition to these theoretical work, an experimental validation of simplified TDIR using the GRACE Follow-On satellites was described in Ref. [27].
In the most recent research regarding Bayesian TDIR [25], the secondary noises, which may drive the estimated parameters away from their true values, are omitted in the simulation. To further investigate the performance of TDIR when secondary noises are also present, in this paper, we apply TDIR on simulated data including the realizations of LFN and secondary noises. For the sake of practicality, the data are generated based on the numerical orbit of Taiji. More importantly, since we aim to develop TDIR into a ranging technique independent of the PRN signal, the implementation of it cannot directly or indirectly rely on the PRN signal. This implies that we need to make some modifications to the algorithm, especially considering its relationship with clock synchronization, as will be explained in the following.
A subtle aspect about TDIR is its relationship with clock synchronization. The laser beatnotes and other measurements are recorded with respect to the clocks (such as ultra-stable oscillators, USOs) hosted by the three S/Cs, each exhibiting a unique deviation from the proper time due to instrumental imperfections including long-term deterministic drifts and random jitters. To relate the data with astrophysical events and in-orbit incidents, thereby enabling the estimation of GW source parameters and monitoring the status of detector, it is essential to adjust the timestamps of data from the three S/Cs to a common time frame (i.e. clock synchronization). In the context of TDI, ranging and synchronization are closely related for at least two reasons. Firstly, during the construction of TDI channels, the time frame of beatnote signals must be consistent with the delays applied on them. For example, if the signals are synchronized to the Solar system barycentric coordinate time (TCB) [28, 29, 30], then the delays must be the light travel time (LTT) defined as the difference of TCB times between photon reception and emission. Ref. [21, 31] predicted that TDI requires this synchronization to reach $\mathcal{O}(10)$-$\mathcal{O}(100)\ {\rm ns}$ precision. While, if the signals are the unsynchronized raw measurements, then the delays should be accordingly the PRNR, defined as the difference of onboard clock readings between photon reception and emission, which encodes the information of both LTT and clock deviations (i.e. the “TDI without synchronization” scheme, see Ref. [32] for more detailed explanations). Secondly, in the conventional data processing pipeline [33], both the nanosecond level high-precision ranging and clock synchronization rely on the PRN signal. An important aspect of PRN signal processing is to disentangle the LTTs and clock deviations using Kalman-like filters, based on the modeling of clocks, orbital motions and relativistic corrections.
Most of the previous studies [23, 24, 25] illustrated TDIR on ideally synchronized data (e.g. synchronization is explicitly stated in Ref. [23], and Ref. [25] modeled the delays as LTTs, hence implicitly assuming synchronized to TCB). As is explained, these implementations of TDIR inherently relied on PRNR. Therefore, using ideally synchronized data contradicts our intention to avoid the dependence of TDIR on the PRN signal. Ref. [22] applied TDIR to unsynchronized raw measurements, while the authors only treated it as an auxiliary step in the so-called “ranging sensor fusion” processing, to correct the bias and ambiguity of PRNR. In this paper, we investigate the feasibility of TDIR on data that has been roughly synchronized to TCB with a $0.1\ {\rm ms}$ precision (still failing to meet the requirement of TDI) using the ground tracking data. As will be explained in Sec. II, in principle, TDIR also applies to completely unsynchronized data. However, for the sake of GW source parameter estimation, synchronization to TCB (or other global time frames) is inevitable. The requirement of GW analysis on clock synchronization is much less stringent than LFN suppression. For the Taiji mission, the specific requirement is under on-going research based on the scientific objectives. In the context of ranging and LFN suppression, this preliminary synchronization offers the opportunity of combining different TDI channels to achieve better constraint on the delay parameters. Accordingly, to encompass the information of both LTT and clock desynchronization into TDIR, we need to adopt a more general parametrization (other than constant delays or Keplerian orbital elements).
Except for LFN, The complete TDI procedure is also known to mitigate the noises induced by the random displacements of optical benches (OBs) relative to the local inertial frames (OB noise), and by the jitters of on board clocks (clock noise). Mitigation of the former noise is accomplished through the construction of test-mass-to-test-mass interferometry (intermediate variable $\xi$), and has been included in previous studies. The latter, however, which should be dealt with an extra clock noise reduction step subsequent to the combination of TDI channels using the information of clock sidebands [34, 35], has received scant attention in the literature about TDIR. In fact, the residual clock noise in the second-generation TDI channel is still 2 - 3 orders larger than the tolerable noise level. If the clock noise reduction step is omitted, whitening the data using only the PSD of secondary noises would be incorrect. Therefore, we suggest that the practical implementation of TDIR should encompass this crucial step. Furthermore, through theoretical analysis and simulation, we investigate the requirements of suppressing OB and clock noises on the precision of TDIR.
In summary, this paper is dedicated to exploring TDIR as an independent ranging technique, under the realistic scenarios including LFN, secondary noises, OB noise, clock noise, orbital motion, and clock desynchronization. The whole TDI procedure based on our TDIR algorithm is free of the imperfections of PRNR, as well as the sophisticated PRNR processing steps, and has more relaxed requirement on clock synchronization. The content of this paper is arranged as follows: Sec. II explains theoretically the principle of applying TDIR to unsynchronized data, and analyzes the couplings of ranging errors with LFN, OB noise and clock noise, thereby deriving the requirements of suppressing these noises on TDIR. Furthermore, the parametrization form of delays is proposed and justified based on the numerical orbit of Taiji. Sec. III then explains how to formulate TDIR as a Bayesian parameter estimation problem, and analytically estimates the accuracy of parameters by means of Fisher matrix. The specific settings for data simulation and MCMC sampling are detailed in Sec. IV, followed by the discussions on the results. We draw our conclusion and outlook for future work in Sec. V. In addition, the models of instrumental noises and an illustrative example regarding the impact of laser locking are presented in Appendix A and B, respectively.
II demonstration of the principles
II.1 TDIR on unsynchronized data
[…]
$\hat{\tau}_{i}^{t}(\tau)=\tau+\delta\hat{\tau}_{i}^{t}(\tau),$ (1)
where the total deviation with respect to TCB is grounded in one term $\delta\hat{\tau}_{i}^{t}(\tau)$. The conversion of $f$ between different time frames thus follows
$f^{\hat{\tau}_{i}}(\tau)=f^{t}\left[t^{\hat{\tau}_{i}(\tau)}\right],$ (2)
Readers can find more transformation rules in Ref. [36].
[…]
$d_{ij}^{t}(\tau)=t_{i,{\rm reception}}-t_{j,{\rm emission}},$ (3)
[…]
$\hat{d}_{ij}^{t}(\tau)$ $=$ $\hat{\tau}_{i}^{t}(\tau)-\hat{\tau}_{j}^{t}[\tau-d^{t}_{ij}(\tau)]$ (4)
$=$ $d^{t}_{ij}(\tau)+\delta\hat{\tau}_{i}^{t}(\tau)-\delta\hat{\tau}_{j}^{t}\left[\tau-d^{t}_{ij}(\tau)\right],$
$\hat{d}^{\hat{\tau}_{i}}_{ij}(\tau)$ $=$ $\hat{d}^{t}_{ij}\left[t^{\hat{\tau}_{i}}(\tau)\right],$ (5)
which incorporates the information of both LTT and differential clock deviations between S/Cs. The first definition is expressed in TCB, and the second gives its transformation to the clock time of receiving S/C, as an application of Eq. (2). To effectively suppress LFN with TDI on unsynchronized data, similar to the “TDI without clock synchronization” scheme [32], the objective of TDIR can be interpreted as searching for the functional form that fits $\hat{d}_{ij}^{\hat{\tau}_{i}}(\tau)$ best. However, perfect reconstruction of $\hat{d}_{ij}^{\hat{\tau}_{i}}(\tau)$ with a fitting function is not practical due to the presence of random clock jitters in $\delta\hat{\tau}_{i}^{t}$. Functional forms with limited number of parameters (as few as possible to ensure the efficiency of MCMC sampling) can at most fit the slowly-varying, deterministic part of $\hat{d}_{ij}^{\hat{\tau}_{i}}(\tau)$. We refer to this variable as the “fiducial” TDIR $\bar{d}_{ij}^{\hat{\tau}_{i}}(\tau)$, to be distinguished it from the “parametrized” TDIR dubbed $R_{ij}(\tau,\bm{\theta})$. The fiducial TDIR will be treated as the “ground truth” of our searching algorithm, and can be calculated via Eq. (5) assuming that there is no random clock jitters. Meanwhile, the parametrized TDIR contains unknown model parameters to be fitted using the beatnote signals.
Next, we explain the principle of suppressing LFN in unsynchronized data using the aforementioned TDIR, starting with the equations of interferometric measurements expressed in terms of TDIR.
[…]
$\textbf{D}_{ij}f^{t}(t)|_{t=t^{\hat{\tau}_{i}}(\tau)}=\textbf{D}_{ij}^{\hat{\tau}_{i}}f^{\hat{\tau}_{j}}(\tau).$ (6)
[…]
$\textbf{D}_{ij}^{\hat{\tau}_{i}}f^{\hat{\tau}_{j}}(\tau)\approx\bar{\textbf{D}}_{ij}^{\hat{\tau}_{i}}f^{\hat{\tau}_{j}}(\tau).$ (7)
In this paper, we construct TDI combinations from detrended data, namely the large phase ramps corresponding to the slow-varying MHz beatnote frequencies have already been removed with methods such as polynomial fitting or bandpass filtering [38, 35], and the only terms left are the ones representing noises and GW signals. The beatnote signals are usually modeled in terms of LTT (e.g. Ref. [18, 34, 24, 25]). Working on the detrended data enables us to recast their expressions in terms of TDIR using Eq. (6) and Eq. (7):
$s_{ij,c}^{\hat{\tau}_{i}}$ $=$ $\bar{\textbf{D}}^{\hat{\tau}_{i}}_{ij}p^{\hat{\tau}_{j}}_{ji}-p^{\hat{\tau}_{i}}_{ij}+\bar{\textbf{D}}^{\hat{\tau}_{i}}_{ij}\Delta_{ji}^{\hat{\tau}_{j}}+\Delta_{ij}^{\hat{\tau}_{i}}-a^{\hat{\tau}_{i}}_{ij,c}q_{i}^{\hat{\tau}_{i}}$ (8)
$+N^{s,c,{\hat{\tau}_{i}}}_{ij},$
$s_{ij,sb}^{\hat{\tau}_{i}}$ $=$ $\bar{\textbf{D}}^{\hat{\tau}_{i}}_{ij}p^{\hat{\tau}_{j}}_{ji}-p^{\hat{\tau}_{i}}_{ij}+\bar{\textbf{D}}^{\hat{\tau}_{i}}_{ij}\Delta_{ji}^{\hat{\tau}_{j}}+\Delta_{ij}^{\hat{\tau}_{i}}-a^{\hat{\tau}_{i}}_{ij,sb}q_{i}^{\hat{\tau}_{i}}$ (9)
$+\nu^{m}_{ji}\bar{\textbf{D}}^{\hat{\tau}_{i}}_{ij}q_{j}^{\hat{\tau}_{j}}-\nu^{m}_{ij}q_{i}^{\hat{\tau}_{i}}+N^{s,sb,{\hat{\tau}_{i}}}_{ij},$
$\tau_{ij,c}^{\hat{\tau}_{i}}$ $=$ $p^{\hat{\tau}_{i}}_{ik}-p^{\hat{\tau}_{i}}_{ij}-b^{\hat{\tau}_{i}}_{ij,c}q_{i}^{\hat{\tau}_{i}}+N^{\tau,c,{\hat{\tau}_{i}}}_{ij},$ (10)
$\varepsilon_{ij,c}^{\hat{\tau}_{i}}$ $=$ $p^{\hat{\tau}_{i}}_{ik}-p^{\hat{\tau}_{i}}_{ij}+2\left(\Delta_{ij}^{\hat{\tau}_{i}}-\delta^{{\hat{\tau}_{i}}}_{ij}\right)-b^{\hat{\tau}_{i}}_{ij,c}q_{i}^{\hat{\tau}_{i}}$ (11)
$+N^{\varepsilon,c,{\hat{\tau}_{i}}}_{ij},$
where $s_{ij}^{\hat{\tau}_{i}},\tau_{ij}^{\hat{\tau}_{i}},\varepsilon_{ij}^{\hat{\tau}_{i}}$ represent the science (inter-spacecraft), reference and test-mass interferometor readouts in the phase unit. Subscript “c”/“sb” is short for the carrier/sideband, while $ij\in\{12,21,23,32,31,13\}$ stands for the index of optical bench where the measurements are taken. On the r.h.s of the equations, $p$, $\Delta$, $\delta$, $q$, $N^{\rm ifo}$ denote the laser phase fluctuation caused by LFN, the phase noises originating from the motions of optical benches and test masses, the clock jitter, and the OMS noise of corresponding interferometer, respectively. Besides, $\nu^{m}_{ij}$ is the 2.4/2.401GHz modulation frequency of sideband, and $a$, $b$ are the MHz beatnote frequencies of corresponding signals, which couple with the clock jitters to form the clock noise. The indices of signals are the same as their corresponding OBs. See FIG. 1 for a schematic layout of Taiji’s constellation.
For the applications of TDI, it is crucial that the delays used to shift the signals must be consistent with the actual delays caused by the physical propagation along the arms (i.e. the delays appearing in the equations of signals). Formally Eq. (8-11) looks similar to their “LTT versions” [18, 34, 24, 25], thus all the deduction and algorithm of TDI still applies, and the only two alterations which should be made are to change the delays from LTT to TDIR, and to interpret the beatnote signals as the unsynchronized measurements. We therefore proved that TDIR can be used to suppress LFN.
By making slight modifications to the standard formulae in the literature, we explicitly present the expressions for the complete TDI data processing flow using TDIR. Firstly, the test-mass-to-test-mass interferometry (denoted as intermediate variable $\xi$) is constructed to eliminate the OB noise:
$\xi_{ij}^{\hat{\tau}_{i}}=s_{ij}^{\hat{\tau}_{i}}+\frac{\tau_{ij}^{\hat{\tau}_{i}}-\varepsilon_{ij}^{\hat{\tau}_{i}}}{2}+\bar{\textbf{D}}_{ij}^{\hat{\tau}_{i}}\frac{\tau_{ji}^{\hat{\tau}_{j}}-\varepsilon_{ji}^{\hat{\tau}_{j}}}{2},$ (12)
with $ij\in\{12,21,23,32,31,13\}$. Secondly, intermediate variable $\eta$ is synthesized to eliminate half of the LFN on each S/C:
$\eta_{ij}^{\hat{\tau}_{i}}=\xi_{ij}^{\hat{\tau}_{i}}+\bar{\textbf{D}}^{\hat{\tau}_{i}}_{ij}\frac{\tau_{ji}^{\hat{\tau}_{j}}-\tau_{jk}^{\hat{\tau}_{j}}}{2},\quad\eta_{ik}^{\hat{\tau}_{i}}=\xi_{ik}^{\hat{\tau}_{i}}+\frac{\tau_{ij}^{\hat{\tau}_{i}}-\tau_{ik}^{\hat{\tau}_{i}}}{2},$ (13)
where $\{i,j,k\}\in\{\{1,2,3\},\{2,3,1\},\{3,1,2\}\}$. The third step is the combination of TDI channels. We take the second-generation Michelson $X_{2}$ channel as an example:
$X_{2}^{\hat{\tau}_{1}}$ $=$ $\left(1-\bar{\textbf{D}}^{\hat{\tau}_{1}}_{121}-\bar{\textbf{D}}^{\hat{\tau}_{1}}_{12131}+\bar{\textbf{D}}^{\hat{\tau}_{1}}_{1312121}\right)\left(\eta_{13}^{\hat{\tau}_{1}}+\bar{\textbf{D}}^{\hat{\tau}_{1}}_{13}\eta_{31}^{\hat{\tau}_{3}}\right)$
-- $\left(1-\bar{\textbf{D}}^{\hat{\tau}_{1}}_{131}-\bar{\textbf{D}}^{\hat{\tau}_{1}}_{13121}+\bar{\textbf{D}}^{\hat{\tau}_{1}}_{1213131}\right)\left(\eta_{12}^{\hat{\tau}_{1}}+\bar{\textbf{D}}^{\hat{\tau}_{1}}_{12}\eta_{21}^{\hat{\tau}_{2}}\right),$
[…]
$X_{2,c}^{\hat{\tau}_{1}}=X_{2}^{\hat{\tau}_{1}}-X_{2,q}^{\hat{\tau}_{1}},$ (15)
$X_{2,q}$ being the clock noise correction term. This step is a necessity for TDIR, or otherwise the TDI data streams would be dominated by the clock noise, making the noise modeling in the TDIR likelihood (using the PSD of secondary noises, see Sec. III) incorrect. Therefore, we suggest that the practical implementation of TDIR should encompass this crucial step. For brevity we will drop the subscript “$c$” in the following, and one should keep in mind that the clock noise reduction step should be added to the TDI data processing flow.
It is worth noting that we have not set any restriction to the physical origin of clock deviation $\delta\hat{\tau}_{i}^{t}(\tau)$. In principle, it can be either the original (uncorrected) clock deviations, or the residual timestamp errors left by a preliminary synchronization using ground tracking data (ground-tracking synchronization hereafter), whose uncertainty is reported to be $\sim 0.1\ {\rm ms}$ [22]. Placing this preliminary synchronization step prior to TDI offers the opportunity to combine different TDI channels, so as to achieve better constraint on the delay parameters.
Besides, it is reported by Ref. [22] that orbital determination provided by ground tracking enables us to calculate LTT [40] (ground-tracking range hereafter) with an accuracy of $50\ {\rm km}$ ($\sim 0.1$ ms). This information can be used in the modeling of secondary noises, and we will demonstrate it allows to accelerate the processing of clock noise in TDIR. We set the ground-tracking ranges to have a 0.1 ms order error relative to LTTs. Our algorithm doesn’t rely on the correlations between this ranging error and the error of ground-tracking synchronization, thus they are assumed independent of each other. The differences between ground-tracking ranges and fiducial TDIR are $\delta\hat{\tau}_{i}^{t}(\tau)-\delta\hat{\tau}_{j}^{t}\left[\tau-d^{t}_{ij}(\tau)\right]$ plus the aforementioned errors of ground-tracking ranges, adding up to total differences of less than 1 ms.
To give a more intuitive explanation of the principles, beatnote signals in the above derivation are all represented in phase units. Following recent progress in the TDI algorithm [41], we perform simulation and TDI processing with data in the frequency unit. Accordingly, all the time delay operators should be replaced by the Doppler delay operators:
$\dot{\bar{\textbf{D}}}_{ij}^{\hat{\tau}_{i}}f=\left(1-\dot{\bar{d}}_{ij}^{\hat{\tau}_{i}}(\tau)\right)\bar{\textbf{D}}_{ij}^{\hat{\tau}_{i}}f.$ (16)
II.2 Requirements of noise suppression on TDIR
Throughout the entire TDI data processing flow, sufficient noise suppression demands the delay operators used in TDI combination to be coincident with the ones appearing in the expressions of signals. The error of TDIR will couple with instrumental noises and result in undesirable noise residuals in the TDI combinations. In this subsection, The coupling of ranging error with LFN, OB noise and clock noise will be investigated through theoretical analysis and simulation. These discussions can be used to set a threshold for the accuracy of TDIR.
The noise cancellation performance of TDI is usually evaluated relative to the secondary noises, mainly including the acceleration noise of test masses, corresponding to the $\delta$ terms in Eq. (8-11), and the OMS noises, corresponding to the $N^{\rm ifo}$ terms. To make an order-of-magnitude analysis, we take the $X_{2}$ channel as an example, and consider the equal-arm case. The PSD of secondary noises in $X_{2}$ reads
$P^{\rm 2nd}_{X_{2}}(f)$ $=$ $64\sin^{2}(2u)\sin^{2}(u)$ (17)
×$\times\left[P_{\rm OMS}(f)+P_{\rm ACC}(f)(\cos^{2}u+3)\right],$
where $u\equiv 2\pi fL/c$, $L=3\times 10^{9}\ {\rm m}$ being the nominal arm-length of Taiji. The instrumental noise models used in the analyses and simulations are detailed in Appendix A. This expression also applies to the $Y_{2}$ an $Z_{2}$ channels. We set 1/$1/10P^{\rm 2nd}_{X_{2}}(f)$ as the thresholds for the noise cancellation capability of TDI.
Next, we investigate the residual noises after TDI caused by the inaccuracy of ranging. For the sake of generality, we do not set a specific laser locking scheme [42, 38] but rather assume all the six lasers are independent. Besides, to derive analytical expressions for a given level of ranging accuracy, the errors of ranging are set as constant biases dubbed $\Delta R_{ij}$. According to Eq. (12-II.1), the PSDs of the residual LFN and OB noise caused by the coupling with ranging errors in the $X_{2}$ channel are
$P^{\rm L}_{X_{2}}(f;\Delta R_{ij})$ $=$ $64\sin^{2}(2u)\sin^{2}(u)\omega^{2}$ (18)
×$\times P_{\rm L}(f)\Delta R^{2}$
$P^{\rm OB}_{X_{2}}(f;\Delta R_{ij})$ $=$ $64\nu_{0}^{2}\sin^{2}(2u)\sin^{2}(u)\omega^{4}$ (19)
×$\times P_{\rm OB}(f)\Delta R^{2},$
where $\omega\equiv 2\pi f$, $\nu_{0}=281.6\ {\rm THz}$ is the central frequency of laser, and
$\Delta R\equiv\sqrt{\frac{\Delta R_{12}^{2}+\Delta R_{21}^{2}+\Delta R_{13}^{2}+\Delta R_{31}^{2}}{4}}$ (20)
defines the average level of ranging errors. Eq. (18) is in consistency with the results of previous studies [32, 43], and one can easily verify that the $\mathcal{O}(10)\ {\rm ns}$ requirement is reasonable for the noise models in consideration. While, under the same $\Delta R$, the residual OB noise is much smaller. These expressions are verified via simulations. The algorithms of clock noise reduction varies in the literature (e.g. [34, 35, 39]), thus the residual clock noise can not be written in a unified analytical form. We adopt the treatment of Ref. [35], and explore the impact of ranging error via a clock-noise-only simulation.
Figure 2: Analytical and simulation results of the residual noises caused by ranging errors in the Michelson $X_{2}$ channel. All the ranging errors are generated from normal distributions $\mathcal{N}(0,\sigma R)$, with the values of $\sigma R$ vary for different noises. The grey, green and orange ASD plots represent the residual LFN, OB noise and clock noise, with the corresponding analytical expressions shown in black dashed line. The red solid and dotted lines stand for 1 and 1/10 times the secondary noises, respectively, which are set as the thresholds for TDI noise suppression.
The analytical and simulation results are visualized in FIG. 2. In our simulation, the ranging errors are generated according to normal distributions $\mathcal{N}(0,\sigma R)$, with the values of $\sigma R$ vary for different noises, which are 10 ns, 10 ns, 1 ms for the LFN, OB noise and clock noise, respectively. The grey, green and orange amplitude spectral density (ASD) plots represent the residual LFN, OB noise and clock noise, with the corresponding analytical expressions shown in black dashed line. The red solid and dotted lines stand for 1 and 1/10 times the secondary noises, respectively, which are set as the thresholds for TDI noise suppression. Besides the consistency between theoretical analysis and simulations, the figure also indicates that although the ground-tracking ranges deviate from the fiducial TDIR for less than 1 ms, they are already sufficient for clock noise correction. This result is beneficial for the practical implementation of TDIR. Currently the post-TDI clock noise algorithms are usually more complex than TDI itself, therefore require more computational time. To tackle this issue, we do not need to compute $X_{2,q}$ at every iteration of the MCMC run, but rather calculate it with the ground-tracking ranges in advance. Furthermore, within the regions of parameter space of our interest (tens of nanoseconds around the fiducial TDIR), OB and clock noises only make negligible contributions to the total noise budget, thus the performance of TDIR is mainly determined by LFN and the secondary noises.
II.3 Parametrization of TDIR
TDIR utilizes parameterized functions as the hypothetical model, and the parameters will be constrained using the beatnote signals. Ref. [25] proposed a fitting form for the delays based on Keplerian orbit. This approach, according to our analysis, is only applicable when the delays are LTTs. Applying TDI to unsynchronized data demands that the delays also include information about clock deviations, hence a more general parametrization is desired. We employ a quadratic polynomial parametrization to fit the delays over a day:
$R_{ij}(\tau;\bm{\theta})=A_{ij}\tau^{2}+B_{ij}\tau+C_{ij},$ (21)
where $\bm{\theta}=\{A_{ij},B_{ij},C_{ij}\},ij\in\{12,21,23,32,31,13\}$ includes 18 parameters in total.
Figure 3: Left panel: the standard deviations of the differences between the fitting and true LTTs (orange curve) / pseudo ranges (green curve), evaluated per day in the whole mission period. Right panel: the ASDs of $X_{2}$ channels constructed from a LFN-only simulation using the fiducial and parametrized TDIRs. All the parameters are determined via least square fit.
The degree of polynomial is closely related to the time duration of data being processed. Due to the complexity of orbital dynamics and relativistic corrections, longer duration naturally requires higher-degree polynomials and more model parameters, hence heightening the challenge of MCMC parameter estimation. On the other hand, as will be analyzed in Sec. III, achieving better constraint on the delay parameters requires the availability of more data points. After balancing between difficulty and precision, we choose the duration of data to be one-day.
Through calculations based on the numerical orbit, we find that quadratic polynomial is sufficient for this timescale. The numerical orbit is provided by Ref. [44, 45], which takes into consideration the influence of the gravitational fields of celestial bodies and is qualified for space-based GW detection. Throughout the whole mission period of 5 years, we compute the LTTs using the formulas in Ref. [37], and fit Eq. (21) within each day using the least square fitting method. The standard deviations of the differences between the fitting and true LTTs are shown with the orange curve in the left panel of FIG. 3. Then, the original (uncorrected) clock deviations are set according to the USO model in Ref. [32], which assumes a $5\times 10^{-7}\ {\rm s/s}$ drift (43 ms/day). We repeat above calculations to the pseudo ranges resulting from this clock model, and the outcome is represented by the green curve in the same panel. For TDIR corresponding to the residual clock deviation left by ground-tracking synchronization, the result lies somewhere between the green and orange curves. According to the threshold established in Sec. II.2, we justify Eq. (21) to be a reasonable parametrization. The right panel of FIG. 3 shows the ASDs of $X_{2}$ channels constructed from a LFN-only simulation using the fiducial and parametrized TDIRs, with the parameters determined via least square fit. It is obviously shown that the result of fiducial TDIR manifests a very ideal LFN suppression effect. The residual LFN resulting from parametrized TDIR is well below 1/10 of the secondary noises, although larger than the fiducial one. The settings underlying this simulation are detailed in Sec. IV.1.
III TDIR in the Bayesian framework
The Bayesian statistical inference framework is a mathematical approach to reasoning about uncertainty, which allows for the incorporation of prior knowledge and noisy data into posterior distributions. According to the Bayes’ theorem, the posterior of model parameter reads
$p\left(\bm{\theta}|d\right)=\frac{\mathcal{L}(d|\bm{\theta})\pi(\bm{\theta})}{p(d)},$ (22)
where $\mathcal{L}(d|\bm{\theta})$ is the likelihood function, namely the distribution of observational data $d$ conditioned on the model parameter $\bm{\theta}$. Without any prior knowledge on the parameters, the prior $\pi(\bm{\theta})$ is set as a multidimensional uniform distribution. The evidence $p(d)$ only acts as a normalization factor and is irrelevant to the values of parameters once the hypothetical model has been chosen, thus
$p(\bm{\theta}|d)\propto\mathcal{L}(d|\bm{\theta}).$ (23)
Assuming that the instrumental noises are Gaussian and stationary, $\mathcal{L}$ can be expressed in the form of matched filtering likelihood:
[…]
$=$ $-2\sum_{f_{k}}\frac{|\tilde{X}(f_{k};\bm{\theta})|^{2}\Delta f}{P^{\rm 2nd}_{X}(f_{k})}=-\sum_{f_{k}}\frac{P_{X}(f_{k};\bm{\theta})}{P^{\rm 2nd}_{X}(f_{k})},$
where $\tilde{X}(f;\bm{\theta})$ represents the Fourier transform of an arbitrary TDI channel named $X$, and $P_{X}(f;\bm{\theta})$ is its one-sided PSD (periodogram). The data is whitten by the theoretical one-sided PSD of the secondary noises $P^{\rm 2nd}_{X}(f)$, since secondary noises dominate the TDI data in the region of parameter space of our interest (tens of nanoseconds around the fiducial TDIR). In the realistic unequal-arm situation, $P^{\rm 2nd}_{X}(f)$ depends on the LTTs, which can take the values of ground-tracking ranges. It can be easily verified that neither the sub-millisecond uncertainties nor the variations of LTTs during a day have considerable impact on $P^{\rm 2nd}_{X}(f)$, using Eq. (40-44) via numerical or analytical calculation. Eq. (24) is equivalent to the ones adopted by previous studies [24, 25]. In the more general “global fit” [46] scenario where GW signal exists, ${\rm ln}\mathcal{L}$ should be replaced by a more familiar form $-1/2(X-h|X-h)$, $h$ being the template of GW signal.
Simultaneously estimating the delay parameters and GW source parameters can be a challenging task. Additionally, Taiji will be able to detect the GWs from tens of millions of compact binaries in the Milky Way, and the majority of them cannot be resolved individually, resulting in a foreground confusion noise in the mHz band [47]. Precise modeling of this noise is only possible when all the bright GW signals have been correctly recognized and subtracted. Therefore, to reduce the encounters with GW signals and foreground, we set $f_{\rm min}=10\ {\rm mHz}$, above the most sensitive band of Taiji. This choice of $f_{\rm min}$ also allows us to accelerate the estimation process by incorporating a preliminary search phase, see Sec. IV. Ideally, under the condition that the computing capacity is sufficient and the algorithm has been optimized, one may discard this phase and raise $f_{\rm min}$ above 0.1 Hz to further minimize the impact of GWs.
In this paper, we perform the time shift of signals with Lagrangian interpolation [31]. Interpolation error increases dramatically near the Nyquist frequency. For a sampling frequency of 4 Hz, we find that Lagrangian interpolation of order 15 can enhance the efficiency of TDI calculation (compared to the widely adopted 31 order) with acceptable compromise in precision. Under this configuration, interpolation error is well below the secondary noises when $f<f_{\rm max}=0.5\ {\rm Hz}$.
The Fisher information matrix is widely employed to predict the uncertainty of parameter estimation in the context of GW observations  [48, 49]. We take the equal-arm case as an example to roughly estimate the precision of the delays using the Fisher matrix formalism. For the simplified analysis, we only consider the Michelson $X_{2}$ channel and assume that the delay for each link is represented by a constant $R_{ij}$. Under these simplifications, the logarithmic likelihood function can be rewritten as
${\rm ln}\mathcal{L}$ $=$ $-\sum_{f_{k}}\frac{P^{\rm L}_{X_{2}}(f_{k};\bm{\theta})+P^{\rm 2nd}_{X_{2}}(f_{k})}{P^{\rm 2nd}_{X_{2}}(f_{k})}$ (25)
$=$ $-\sum_{f_{k}}\left[1+\frac{P^{\rm L}_{X_{2}}(f_{k};\bm{\theta})}{P^{\rm 2nd}_{X_{2}}(f_{k})}\right].$
When using this single channel, it can be easily verified that there is no correlation between different arms via definition
${\rm Cov}_{\theta_{a}\theta_{b}}\equiv\frac{\Gamma^{-1}_{\theta_{a}\theta_{b}}}{\sqrt{\Gamma^{-1}_{\theta_{a}\theta_{a}}\Gamma^{-1}_{\theta_{b}\theta_{b}}}},$ (26)
where
$\Gamma_{\theta_{a}\theta_{b}}=-\frac{\partial^{2}{\rm ln}\mathcal{L}}{\partial\theta_{a}\partial\theta_{b}}.$ (27)
is the element of Fisher matrix corresponding to parameters $\theta_{a}$ and $\theta_{b}$, and $\Gamma^{-1}$ denotes the inverse matrix of $\Gamma$. Therefore, the Cramer-Rao bounds for the uncertainties of delays are
$\sigma_{R_{ij}}^{-2}$ $=$ $\Gamma_{R_{ij}R_{ij}}=-\frac{\partial^{2}{\rm ln}\mathcal{L}}{\partial R_{ij}^{2}}=\sum_{f_{k}}\frac{\frac{\partial^{2}P^{\rm L}_{X_{2}}}{\partial R_{ij}^{2}}(f_{k},\bm{\theta})}{P^{\rm 2nd}_{X_{2}}(f_{k})}$ (28)
$\approx$ $\frac{c^{2}}{2\nu_{0}^{2}}\frac{A_{\rm L}^{2}}{A_{\rm OMS}^{2}}N,\quad ij\in\{12,21,13,31\}.$
The last equivalence holds due to the fact that OMS noise dominates the noise budget at the frequency range of interest. Therefore
$\sigma_{R_{ij}}\approx\frac{\sqrt{2}\nu_{0}}{c}\frac{A_{\rm OMS}}{A_{\rm L}}\frac{1}{\sqrt{N}}\approx\frac{3.54\times 10^{-7}\ {\rm s}}{\sqrt{N}}.$ (29)
For $f_{\rm min}=10\ {\rm mHz}$, $f_{\rm max}=0.5\ {\rm Hz}$ and $df=1/{\rm day}$, the number of data points in the Fourier domain is $N=42336$, thus $\sigma_{R_{ij}}\approx 1.72\ {\rm ns}$. This represents a rather optimistic estimate, as more complex parametrizations would likely introduce correlations between the parameters characterizing each arm, which could inflate the errors.
Figure 4: The ASDs of secondary noises and residual LFN in the $A_{2}$ and $T_{2}$ channels, with the armlengths calculated from the orbit data in the 1st day, and a 10 ns level ranging error. The shapes of $E_{2}$ channel are similar to those of $A_{2}$ hence not shown.
The data of a single TDI channel $X_{2}$ can only place constraints on $R_{12}$, $R_{21}$, $R_{13}$ and $R_{31}$. Using multiple channels is necessary for sampling the entire parameter space, and it can also help to obtain tighter constraints on the parameters. In further analysis based on the simulated data, to fully utilize the information from different TDI channels, and to maintain consistency with the conventions of GW source parameter estimation so as to be integrated into the global fit pipeline, we employ the “optimal” TDI channels constructed from the Michelson channels as [50]:
$A_{2}$ $=$ $\frac{Z_{2}-X_{2}}{\sqrt{2}},$ (30)
$E_{2}$ $=$ $\frac{X_{2}-2Y_{2}+Z_{2}}{\sqrt{6}},$ (31)
$T_{2}$ $=$ $\frac{X_{2}+Y_{2}+Z_{2}}{\sqrt{3}},$ (32)
and the total logarithmic likelihood function is the sum of $\{A_{2},E_{2},T_{2}\}$ channels. The PSDs of secondary noises in the optimal channels can be found in Appendix A, where we have also presented the PSDs of residual LFN due to the coupling with ranging errors. Observing Eq. (25), the contribution of a certain TDI channel to the total likelihood is determined by the ratio between residual LFN and secondary noises. These two terms for the $\{A_{2},T_{2}\}$ channels are shown in FIG. 4, with the armlengths calculated from the numerical orbit data in the 1st day, and a 10 ns level ranging error. The shapes of noises in $E_{2}$ are similar to those of $A_{2}$ hence not shown. As can be seen from the figure, within the frequency band in concern, each channel contributes similarly to the likelihood.
[…]
IV Simulation results
IV.1 Settings for data simulation and MCMC sampling
Until now, the functional forms of residual clock errors left by ground-tracking synchronization for the Taiji mission are still under investigation. We assume their form as
$\delta\hat{\tau}_{i}^{t}(\tau)=\sum_{m=0}^{2}\mathcal{C}_{m,i}\tau^{m},\quad i\in\{1,2,3\},$ (33)
with the coefficients generated from normal distributions $\mathcal{C}_{m,i}\sim\mathcal{N}(0,0.1\ {\rm ms}/T_{\rm obs}^{m})$, $T_{\rm obs}=1\ {\rm day}$. Each term contributes 0.1 ms to the total deviation. According to the theoretical analysis in Sec. II, we do not expect the specific shape of $\delta\hat{\tau}_{i}^{t}(\tau)$ to have significant impact on the feasibility of TDIR. Regarding parametrization Eq. (21), the left panel of FIG. 3 shows that it is even applicable to the raw clock deviations, and the combination of different channels is valid up to the limits set at the end of Sec. III. Besides, we assume the ground-tracking range in each of the six laser links to deviate from corresponding LTT following the same form as Eq. (33). Note that these coefficients are independent of $\mathcal{C}_{m,i}$. These ranges are used as LTTs to calculate the theoretical PSDs in the likelihood. The simulation of unsynchronized beatnote signals in the frequency unit follows the approach described in Ref. [36, 38]. To be specific, the signals are initially generated in the proper time of each S/C, then transferred to TCB, and finally resampled according to the clock deviation model set by Eq. (33). To set a benchmark, we calculate the fiducial TDIR according to Eq. (4-5) and obtain the “true values” of parameters via least square fitting.
We generate two sets of simulated data:
  1. (1)
    Simulation I: only contains the realization of LFN (labeled as “only laser frequency noise” in the figures);
  2. (2)
    Simulation II: contains the realizations of LFN as well as secondary noises (labelled as “with secondary noises” in the figures).
The first simulation is intended to validate our parametrization and TDIR algorithm in an ideal situation, while the second one aims to test the performance of the algorithm under a more realistic scenario. The simulations are all based on the 1st day of the numerical orbit, and the sampling frequency is set as 4 Hz. All the instrumental noises are assumed to be Gaussian and stationary, and see Appendix A for the noise models in the form of PSD. Still, for generality no specific laser locking scheme is set and all six lasers are independent. Clock noise and OB noise are not present in either of the simulations since they do not contribute considerably to the likelihood, and the clock noise correction term can be calculated prior to TDIR, as is explained in Sec. II.2.
We use the affine-invariant MCMC ensemble sampler emcee to draw samples from the likelihood [52]. The stretch parameter is set to the default value $a=2$. 72 walkers (4 times the number of parameters) are employed to fully explore the shape of posterior. The whole estimation process can be divided into two phases. The first phase is a preliminary search, where we downsample the data to 1 Hz, and the interpolation order is set to 7, all for the sake of efficiency. $f_{\rm max}$ is correspondingly reduced to 20 mHz to avoid interpolation error. The preliminary search starts from 1 ms intervals centered at the ground-tracking ranges. Once the preliminary search reaches a steady posterior distribution, the second phase, namely the refined search begins from the resultant $3\ \sigma$ confidence intervals. This phase is based on the original 4 Hz data, and we adopt the settings given in Sec. III, i.e. interpolation order 15 and $f_{\rm max}=0.5\ {\rm Hz}$. For the refined search, we run at least 50 auto-correlation times after burn-in following the instruction of the code.
IV.2 Results of MCMC sampling
Figure 5: Posterior distributions of the TDIR parameters. Results based on the LFN-only and the LFN + secondary noises are shown in green and orange corner plots, respectively, where the inner and outer contours represent the 1 $\sigma$ and 2 $\sigma$ ranges of the parameters, and the “true values” are marked with the red lines. In the 1D distribution plots on the diagonal, the vertical dashed lines indicate the 1 $\sigma$ ranges and median values of the corresponding parameter.
Figure 6: Left panel: The distributions of $\Delta R$ calculated from the MCMC samples of Simulation I, II. $\Delta R_{\rm median,I}=0.49\ {\rm ns}$ and $\Delta R_{\rm median,II}=5.28\ {\rm ns}$ are marked with vertical dashed lines. Right panel: The LFN reduction effect achieved by the median values of parameters. Results shown here are also based on a LFN-only simulation, and the ASD in grey indicates the original readout of science interferometer 12 (detrended). The black dashed lines are the theoretical ASDs of residual LFN calculated with $\Delta R_{\rm median,I/II}$.
The posterior distributions obtained from the two simulations are shown in FIG. 5. Between the two overlapping corner plots, results of Simulation I and II are shown in green and orange, respectively. The inner and outer contours represent the 1 $\sigma$ and 2 $\sigma$ confidence intervals of the parameters, and the “true values” are marked with the red lines. In the 1D distribution plots on the diagonal, the vertical dashed lines indicate the median values and 1 $\sigma$ ranges of corresponding parameters. Unsurprisingly, the posterior distribution based on Simulation I is unbiased, since in the absence of the random disturbance caused by secondary noises, parameters that lead to the maximum ${\rm ln}\mathcal{L}$) (i.e. minimum noise) should coincide with the true values. As for Simulation II, the median values of parameters are shifted relative to the true values, but generally still within the 3 $\sigma$ ranges. Through the comparison between two simulations, we can justify that these shifts are not caused by improper parametrization, but rather due to the impact of the specific realization of secondary noises. The strong degeneracy among the polynomial coefficients describing each arm (e.g. $\{A_{12},B_{12},C_{12}\}$) is a reasonable outcome, as the delays are represented by their combinations. While, no evident correlation between the arms is observed, coincident with the theoretical analysis using a single $X_{2}$ channel.
How to evaluate the accuracy of ranging based on the posterior distribution can be a perplexing issue. This is because the formulas and threshold presented in Sec. II.2 are only suitable for constant biases, whereas we model the delays as functions of time. The deviations of $C_{ij}$ relative to the true values can be up to 100 ns, which seems to suggest a pessimistic conclusion. However, these parameters only represent the initial delays and does not encompass how the fitting functions performs over the whole duration of a day. Therefore, in order to equivalent the time-varying delays to constant biases, hence gaining an intuitive understanding of the results, we define the average ranging error based on the effect of LFN suppression following Eq. (18) and Eq. (20):
$\Delta R\equiv\sqrt{\frac{1}{3N}\sum_{X\in\{X_{2},Y_{2},Z_{2}\}}\sum_{i}^{N}\frac{{P}^{\rm L,MCMC}_{X}(f_{k})}{P^{\rm L}_{X}(f_{k};\Delta R=1)}}.$ (34)
For each MCMC sample, we calculate the corresponding delays with Eq. (21). The delays are then applied to a LFN-only simulated data (not necessarily Simulation I) to combine TDI channels $\{X_{2},Y_{2},Z_{2}\}$, and ${P}^{\rm L,MCMC}_{X}(f_{k})$ is the resultant PSD of channel $X$ at frequency $f_{k}$ Ratio ${P}^{\rm L,MCMC}_{X}(f_{k})/P^{\rm L}_{X}(f_{k};\Delta R=1)$ is then averaged across frequencies $10^{-3}\ {\rm Hz}<f_{k}<2.4\times 10^{-2}\ {\rm Hz}$ (we set this upper limit to avoid modeling errors at the characteristic frequencies) and TDI channels $\{X_{2},Y_{2},Z_{2}\}$ to construct an average measurement of the ranging bias. We have chosen the original Michelson channels instead of the optimal ones since they offer an intuitive way to define an averaged bias $\Delta R$.
The left panel of FIG. 6 shows the distributions of $\Delta R$ calculated from the MCMC samples of Simulation I and II. The $\Delta R$s corresponding to the median values of parameters are represented by the vertical dashed lines and denoted as $\Delta R_{\rm median,I/II}$. For Simulation I, $\Delta R_{\rm median,I}=0.49\ {\rm ns}$ appears at the minimum of the distribution. While, for Simulation II, due to the disturbance cause by the secondary noises, $\Delta R_{\rm median,II}=5.28\ {\rm ns}$ does not coincide with the best LFN suppression effect among the posterior samples, but rather occurs in the lower-middle part of the distribution. Additionally, each distribution of $\Delta R$ exhibit a dispersion of around 2 ns, which, in a sense, can be regarded as a measure for the uncertainty of TDIR.
The LFN reduction effects achieved by the median values of parameters are shown in the right panel of FIG. 6. These results are also based on a LFN-only simulation, and the grey ASD plot indicates the original readout of science interferometer 12. The black dashed lines are the theoretical ASDs of residual LFN calculated with $\Delta R_{\rm median,I/II}$. The agreement between theory and simulation suggests that our definition of the average ranging bias is reasonable. Judging from the threshold set by 1/10 of the secondary noises, we draw a conclusion that TDIR remains feasible in the presence of secondary noises.
In all the aforementioned investigation, for the sake of generality, laser locking was not implemented hence all six lasers are assumed independent of each other. While, in practice, once a specific locking scheme is set, the six lasers become correlated, which would consequently affect the residual LFN after TDI, and the covariances between different arms obtained by TDIR. In Appendix B, we take the locking scheme N1-LA12 [38] as an example to offer a glimpse at the posterior distribution of delay parameters in the presence of laser locking, compared to the “no locking” case. Different correlations between the arms are clearly manifested in FIG. 7. The differences introduced by various locking schemes, as well as the selection of “optimal” scheme in the sense of TDIR remains to be comprehensively explored in future research.
V Conclusions
In this paper, we explored the feasibility and performance of TDIR as a ranging technique independent of the PRN signal, considering the realistic scenarios including LFN, secondary noises, OB noise, clock noise, orbital motion, and clock desynchronization. MCMC sampling based on simulated data yielded a 5.28 ns ranging bias together with a 2 ns dispersion, lower than the threshold which is set according to the effect of LFN suppression.
Our research has shown promising prospects for the application of TDIR, which may serve as a cross-validation or backup for the traditional PRN-based approach. This approach is free of the imperfections of PRNR and the sophisticated PRNR processing steps. Additionally, the whole TDI data processing flow based on our TDIR algorithm has more relaxed requirement on the accuracy of clock synchronization.
Meanwhile, there are at least three issues that remain to be addressed in the future. Firstly, in our research, running the two-phase MCMC sampling took around $\mathcal{O}(5)$ days on a personal laptop, highlighting the imperative to enhance both the computational power and the efficiency of algorithm. Secondly, since TDIR is driven by the interferometric measurements, the components of data have a significant impact on the outcome. In the presence of GW signals, it is necessary to incorporate TDIR into the global fit pipeline, and the effect of joint estimation with GW source parameters necessitates further investigation. Thirdly, the majority of our research assumed six independent lasers, and a simple illustrative example was provided for the case with laser locking. More detailed discussion regarding the impacts and choice of different locking schemes also awaits future research.
Acknowledgements.
This study is supported by the National Key Research and Development Program of China (Grant No. 2021YFC2201903, Grant No. 2021YFC2201901 and Grant No. 2020YFC2200100).
Appendix A Noise models
All instrumental noises are assumed to be Gaussian and stationary, and the amplitudes and shapes of PSDs are chosen according to the current requirements for the Taiji mission.
  • LFN:
    $P_{\rm L}(f)=A_{\rm L}^{2},$ (35)
    with $A_{\rm L}=30{\rm Hz/\sqrt{Hz}}$.
  • OB noise:
    $P_{\rm OB}(f)=A_{\rm OB}^{2},$ (36)
    with $A_{\rm OB}=5.65\times 10^{-18}{\rm s/\sqrt{Hz}}$, equivalent to a nanometer level displacement jitter, or a 10 mrad level phase disturbance.
  • clock noise:
    $P_{\rm clock}(f)=A_{\rm clock}^{2}\left(\frac{{\rm 1mHz}}{f}\right)^{3},$ (37)
    where $A_{\rm clock}=3.16\times 10^{-10}\ {\rm s/\sqrt{Hz}}$.
  • secondary noises:
    $P_{\rm OMS}(f)$ $=$ $A_{\rm OMS}^{2}\left[1+\left(\frac{2{\rm mHz}}{f}\right)^{4}\right],$ (38)
    $P_{\rm ACC}(f)$ $=$ $A_{\rm ACC}^{2}\left[1+\left(\frac{0.4{\rm mHz}}{f}\right)^{2}\right]\left[1+\left(\frac{f}{8{\rm mHz}}\right)^{4}\right].$ (39)
    To achieve the target sensitivity of Taiji, the required amplitudes of optical measurement noise and test-mass acceleration noise are $A_{\rm OMS}=8\ {\rm pm/\sqrt{Hz}}$ and $A_{\rm ACC}=3\ {\rm fm/s^{2}/\sqrt{Hz}}$, respectively. Note that $P_{\rm OMS}(f)$ and $P_{\rm ACC}(f)$ should be multiplied by $(2\pi f\nu_{0}/c)^{2}$ and $(\nu_{0}/(2\pi fc))^{2}$ to be converted to frequency unit before inserting into Eq. (17).
The PSDs of secondary noises in the Michelson and optimal TDI channels for the general unequal-arm scenario read
$P^{\rm 2nd}_{X_{2}}(f)$ $=$ $32\sin^{2}\left(\frac{u_{12131}}{2}\right)\left\{P_{\rm OMS}(f)\left[\sin^{2}\left(\frac{u_{121}}{2}\right)+\sin^{2}\left(\frac{u_{131}}{2}\right)\right]\right.$ (40)
$+2P_{\rm ACC}(f)\left[\sin^{2}\left(\frac{u_{121}}{2}\right)\left(\cos^{2}\left(\frac{u_{131}}{2}\right)+1\right)\right.+\left.\left.\sin^{2}\left(\frac{u_{131}}{2}\right)\left(\cos^{2}\left(\frac{u_{121}}{2}\right)+1\right)\right]\right\},$
$P_{X_{2}Y_{2}^{*}}^{\rm 2nd}(f)$ $=$ $-(1-\Delta_{131})(1-\Delta_{12131})(1-\Delta_{232}^{*})(1-\Delta_{23212}^{*})(\Delta_{21}^{*}+\Delta_{12})\left[P_{\rm OMS}(f)+4P_{\rm ACC}(f)\right],$ (41)
$P_{A_{2}}^{\rm 2nd}(f)$ $=$ $\frac{1}{2}\left[P^{\rm 2nd}_{X_{2}}(f)+P^{\rm 2nd}_{Z_{2}}(f)\right]-{\rm Re}\left[P_{Z_{2}X_{2}^{*}}^{\rm 2nd}(f)\right],$ (42)
$P_{E_{2}}^{\rm 2nd}(f)$ $=$ $\frac{1}{6}\left\{P^{\rm 2nd}_{X_{2}}(f)+4P^{\rm 2nd}_{Y_{2}}(f)+P^{\rm 2nd}_{Z_{2}}(f)+{\rm Re}\left[2P_{Z_{2}X_{2}^{*}}^{\rm 2nd}(f)-4P_{X_{2}Y_{2}^{*}}^{\rm 2nd}(f)-4P_{Y_{2}Z_{2}^{*}}^{\rm 2nd}(f)\right]\right\},$ (43)
$P_{T_{2}}^{\rm 2nd}(f)$ $=$ $\frac{1}{3}\left\{P^{\rm 2nd}_{X_{2}}(f)+P^{\rm 2nd}_{Y_{2}}(f)+P^{\rm 2nd}_{Z_{2}}(f)+2{\rm Re}\left[P_{Z_{2}X_{2}^{*}}^{\rm 2nd}(f)+P_{X_{2}Y_{2}^{*}}^{\rm 2nd}(f)+P_{Y_{2}Z_{2}^{*}}^{\rm 2nd}(f)\right]\right\},$ (44)
where $u_{i_{1}...i_{n}}\equiv 2\pi f(d_{i_{1}i_{2}}+...+d_{i_{n-1}i_{n}})$, $\Delta_{i_{1}...i_{n}}\equiv\exp(-iu_{i_{1}...i_{n}})$, $d_{i_{1}i2},...,d_{i_{n-1}i_{n}}$ being the LTTs. The expressions for $Y_{2}$, $Z_{2}$, $Y_{2}Z_{2}^{*}$ and $Z_{2}X_{2}^{*}$ can be obtained from Eq. (40) and Eq. (41) via cyclical permutation of the indices : $1\rightarrow 2$, $2\rightarrow 3$, $3\rightarrow 1$.
We also present the PSDs of residual LFN in the optimal channels due to ranging errors, as a complement to the $X_{2}$ channel discussed in the main text:
$P^{\rm L}_{A_{2}}(f;\Delta R_{ij})$ $=$ $8\sin^{2}(2u)\sin^{2}(u)\omega^{2}P_{\rm L}(f)\left[\Delta R_{12}^{2}+\Delta R_{21}^{2}+\Delta R_{23}^{2}+\Delta R_{32}^{2}+4\cos^{2}\frac{u}{2}(\Delta R_{31}^{2}+\Delta R_{13}^{2})\right],$ (45)
$P^{\rm L}_{E_{2}}(f;\Delta R_{ij})$ $=$ $\frac{8}{3}\sin^{2}(2u)\sin^{2}(u)\omega^{2}P_{\rm L}(f)$ (46)
×$\times\left[4\sin^{2}\frac{u}{2}(\Delta R_{13}^{2}+\Delta R_{31}^{2})+(5+4\cos u)(\Delta R_{23}^{2}+\Delta R_{32}^{2}+\Delta R_{12}^{2}+\Delta R_{21}^{2})\right],$
$P^{\rm L}_{T_{2}}(f;\Delta R_{ij})$ $=$ $\frac{64}{3}\sin^{2}(2u)\sin^{2}(u)\omega^{2}P_{\rm L}(f)\sin^{2}\frac{u}{2}\left(\Delta R_{12}^{2}+\Delta R_{21}^{2}+\Delta R_{13}^{2}+\Delta R_{31}^{2}+\Delta R_{23}^{2}+\Delta R_{32}^{2}\right).$ (47)
These expressions are derived under the equal-arm approximation and only used in order-of-magnitude analysis.
Appendix B An illustrative example for the case with laser locking
Figure 7: Similar to FIG. 5, but the green and orange corner plots represent the posteriors of the “no locking scheme” and “N1-LA12 locking scheme” cases, respectively. Both results are obtained based on LFN-only simulations.